OpenStudy (anonymous):

wht is 5p4 the 5 and 4 r small

OpenStudy (phi):

you have to learn the definition. 5P4 means \[ \frac{5!}{4!} \] where 5! means 5 factorial , 5!= 5*4*3*2*1

OpenStudy (phi):

if you expand out the factorials \[ 5P4= \frac{5\cdot 4\cdot 3\cdot 2\cdot 1}{4\cdot 3\cdot 2\cdot 1} \] notice the 4/4 and 3/3 etc all divide out (become 1) so this simplifies to 5

OpenStudy (anonymous):

the answer is 5 ??

OpenStudy (arindameducationusc):

yes answer is 5, good job